# Linear Transformers are Versatile In-Context Learners

### v1

@@ Appendix A Omitted proofs from Section 4
By Lemma B.2 we know in $O(\log\kappa+\log\log 1/\epsilon)$ iterations, by assigning $\kappa$ in the way of Lemma B.2 one can reduce the condition number of $x$ to $\kappa^{\prime}\leq 1+\epsilon/2\kappa$ (we chose $\epsilon/2\kappa$ here to give some slack for later analysis).
Let $\Sigma^{\prime}$ be the covariance matrix after these iterations, and $\nu^{\prime},\mu^{\prime}$ be the upper and lowerbound for its eigenvalues. The data $x_{i}$’s are transformed to a new data $x_{i}^{\prime}=Mx_{i}$ for some matrix $M$. Let $M=A\Sigma^{-1/2}$, then since $M^{\prime}=AA^{\top}$ we know $A$ is a matrix with singular values between $\sqrt{\mu^{\prime}}$ and $\sqrt{\nu^{\prime}}$. The optimal solution $(w^{*})^{\prime}=M^{-\top}w^{*}$ has norm at most $\sqrt{\nu}/\sqrt{\mu^{\prime}}\|w^{*}\|$. Therefore by Lemma B.1 we know the one-step gradient step with $\hat{w}=\sum_{i=1}^{n}\frac{1}{\mu^{\prime}}y_{i}x_{i}$ satisfy $\|\hat{w}-(w^{*})^{\prime}\|\leq(\kappa^{\prime}-1)\sqrt{\nu}/\sqrt{\mu^{\prime}}\|w^{*}\|$. The test data $x_{t}$ is also transformed to $x_{t}^{\prime}=A\Sigma^{-1/2}x_{t}$, and the algorithm outputs $\langle\hat{w},x_{t}^{\prime}\rangle$, so the error is at most $\sqrt{\nu}\|w^{*}\|*\|x_{t}^{\prime}\|\leq(\kappa^{\prime}-1)\sqrt{\kappa}\sqrt{\kappa^{\prime}}\|w^{*}\|*\|x_{t}\|$. By the choice of $\kappa^{\prime}$ we can check that RHS is at most $\epsilon\|w^{*}\|\|x_{t}\|$. ∎
Appendix C More experiments
@@
Fig. 6 shows per-variance profile of intermediate predictions of the network of varying depth. It appears that ${}^{++}$ demonstrates behavior typical of GD-based algorithms: early iterations model higher noise (similar to early stopping), gradually converging towards lower noise predictions. Diag exhibits this patter initially, but then dramatically improves, particularly for lower noise ranges. Intriguingly, Full displays the opposite trend, first improving low-noise predictions, followed by a decline in higher noise prediction accuracy, especially in the last layer.
@@
Table 1: Adjusted evaluation loss for models with various number of layers with uniform noise variance $\sigma_{\tau}\sim U(0,\sigma_{max})$. We highlight in bold the best results for each problem setup (i.e. column).

### v2

@@ Appendix A Proofs from Sections 4 and  5
By Lemma A.3 we know in $O(\log\kappa+\log\log 1/\epsilon)$ iterations, by assigning $\kappa$ in the way of Lemma A.3 one can reduce the condition number of $x$ to $\kappa^{\prime}\leq 1+\epsilon/2\kappa$ (we chose $\epsilon/2\kappa$ here to give some slack for later analysis).
Let $\Sigma^{\prime}$ be the covariance matrix after these iterations, and $\nu^{\prime},\mu^{\prime}$ be the upper and lowerbound for its eigenvalues. The data $x_{i}$’s are transformed to a new data $x_{i}^{\prime}=Mx_{i}$ for some matrix $M$. Let $M=A\Sigma^{-1/2}$, then since $M^{\prime}=AA^{\top}$ we know $A$ is a matrix with singular values between $\sqrt{\mu^{\prime}}$ and $\sqrt{\nu^{\prime}}$. The optimal solution $(w^{*})^{\prime}=M^{-\top}w^{*}$ has norm at most $\sqrt{\nu}/\sqrt{\mu^{\prime}}\|w^{*}\|$. Therefore by Lemma A.2 we know the one-step gradient step with $\hat{w}=\sum_{i=1}^{n}\frac{1}{\mu^{\prime}}y_{i}x_{i}$ satisfy $\|\hat{w}-(w^{*})^{\prime}\|\leq(\kappa^{\prime}-1)\sqrt{\nu}/\sqrt{\mu^{\prime}}\|w^{*}\|$. The test data $x_{t}$ is also transformed to $x_{t}^{\prime}=A\Sigma^{-1/2}x_{t}$, and the algorithm outputs $\langle\hat{w},x_{t}^{\prime}\rangle$, so the error is at most $\sqrt{\nu}\|w^{*}\|*\|x_{t}^{\prime}\|\leq(\kappa^{\prime}-1)\sqrt{\kappa}\sqrt{\kappa^{\prime}}\|w^{*}\|*\|x_{t}\|$. By the choice of $\kappa^{\prime}$ we can check that RHS is at most $\epsilon\|w^{*}\|\|x_{t}\|$. ∎
A.5 Proof of the Theorem 5.2
Proof.
The key observation here is that when $n\to\infty$, under the assumptions we have $\lim_{n\to\infty}\sum_{i=1}^{n}x_{i}x_{i}^{\top}=I$. Therefore the ridge regression solutions converge to $w^{*}_{\sigma^{2}}=\frac{1}{1+\sigma^{2}}\sum_{i=1}^{n}y_{i}x_{i}$ and the desired output is $\langle w^{*}_{\sigma^{2}},x_{q}\rangle$.
By the calculations before, we know after the first-layer, the implicit $w$ is $w^{1}=\omega_{yx}\sum_{i=1}^{n}y_{i}x_{i}$. As long as $\omega_{yx}$ is a constant, when $n\to\infty$ we know $\frac{1}{n}\sum_{i=1}^{n}(y^{1}_{i})^{2}=\sigma^{2}$ (as the part of $y$ that depend on $x$ is negligible compared to noise), therefore the output of the second layer satisfies
$w^{2}=(1+n\sigma^{2}\omega_{yy})w^{1}=(1+n\sigma^{2}\omega_{yy})\omega_{yx}\sum_{i=1}^{n}y_{i}x_{i}.$
Therefore, as long as we choose $\omega_{yx}$ and $\omega_{yy}$ to satisfy $(1+n\sigma^{2}\omega_{yy})\omega_{yx}=\frac{1}{1+\sigma^{2}}$ when $\sigma=\sigma_{1}$ or $\sigma_{2}$ (notice that these are two linear equations on $\omega_{yx}$ and $n\omega_{yx}\omega_{yy}$, so they always have a solution), then we have $\lim_{n\to\infty}w^{2}=w^{*}_{\sigma^{2}}$ for the two noise levels. ∎
Appendix B More experiments
@@
Fig. 7 shows per-variance profile of intermediate predictions of the network of varying depth. It appears that GD++ demonstrates behavior typical of GD-based algorithms: early iterations model higher noise (similar to early stopping), gradually converging towards lower noise predictions. Diag exhibits this patter initially, but then dramatically improves, particularly for lower noise ranges. Intriguingly, Full displays the opposite trend, first improving low-noise predictions, followed by a decline in higher noise prediction accuracy, especially in the last layer.
@@
Table 1: Adjusted evaluation loss for models with various number of layers with uniform noise variance $\sigma_{\tau}\sim U(0,\sigma_{max})$. We highlight in bold the best results for each problem setup (i.e. each column).